Use the method of disks/rings to determine the volume of the solid obtained by rotating the region bounded by y=√x, y=3 and the y-axis about the y-axis.


1
Expert's answer
2021-04-15T07:40:17-0400

Here the region is rotating about y-axis as shown below



Volume of the solid obtained by

rotating the region bounded by y=√x, y=3 and the y-axis about the y-axis.

= "\int_{0}^{3} πx²dy"

= "π\int_{0}^{3} x²dy"

= "π\int_{0}^{3} y ^4dy" since y = √x

= "π[ \frac{y^5}{5}]_{0}^{3}"


= π("\frac{3^5}{5}" )


= "\frac{243π}{5}"

= 48.6π

So the required volume will be "\frac{243π}{5}" cubic unit = 48.6π cubic unit